a solution is divided into two smaller solutions P and Q .What volume of 1Molar P and 0.1 Molar Q is required to prepare 2 litres of 0.1 molar solution ?
Answer by checkley79(3341)

1X+.1(2-X)=.1*2
X+.2-.1X=.2
.9X=2-2
.9X=0
X=0 NO MOLAR P IS USED.
PROOF:
0+.1(2-0)=.1*2
0+.1*2=.2
.2=.2